OpenStudy (anonymous):

Solve: 2x^2+12x+9=0 So far I can't get any numbers that multiply to be 18 and add to be 12...

OpenStudy (mathstudent55):

There aren't any two such numbers. That's a good observation on your part. That means you need to use the quadratic formula since the trinomial cannot be factored.

OpenStudy (anonymous):

Thanks, is there an informal rule that is loosely used to tell people that if you can't factor an equation, use the quadratic formula?

OpenStudy (mathstudent55):

Yes. A great question. Use the discriminant of the quadratic formula: \( b^2 - 4ac\) You can only factor the trinomial if the discriminant is a perfect square. Let's l;ook at your problem. \(b^2 - 4ac\) for a = 2, b =- 12, c = 9 gives us: \(12^2 - 4(2)(9) \) \(= 144 - 72\) \(= 72\) 72 is not a perfect square. Its square root simplifies to: \( 3\sqrt{2} \)